OpenStudy (anonymous):

4.The volume of a right circle cone is 800 mm³ and its radius is 5 mm. Find the height of the cone. Use 3.14 for π.

mathslover (mathslover):

Do you know the formula of a right circular cone Hey guys please dont post direct solution ..

OpenStudy (anonymous):

right circular cone volume =1/3pi r^2h=800 3.14*5*5*h=2400 h=30.5

OpenStudy (anonymous):

height of the cone 30.5mm

mathslover (mathslover):

Formula : Volume of a right circular cone = \(\frac{1}{3}\pi r^2h\) given radius = 5 mm

mathslover (mathslover):

and volume = 800 \(mm^3\)

mathslover (mathslover):

Put the values. .. . . may i know wht r u getting